# Manuel Faysse1,5  Patrick Fernandes6,8,11  Nuno M. Guerreiro2,5,6,8 António Loison1  Duarte M. Alves6,8  Caio Corro9  Nicolas Boizard4,5 João Alves2  Ricardo Rei2,7,8  Pedro H. Martins2  Antoni Bigata Casademunt10

### v4

@@ 1 Introduction
Large Language Models (LLM) have taken over the Natural Language Processing (NLP) landscape in the past years. Although a few proprietary models are still considered to run ahead of the pack (OpenAI et al., 2023), open weights models such as Llama (Touvron et al., 2023a, b), Qwen (Bai et al., 2023a) or Mistral (Jiang et al., 2023, 2024) are rapidly bridging the performance gap. However, widespread industrial and research adoption of such technology remains challenging for several reasons, including the lack of transparency in the data collection and training processes, the scarcity of existing resources outside of English, and the large-scale and costly nature of existing high-performing models.
@@
Challenging to use at scale. Although benefits of scaling models to enormous sizes have been amply demonstrated in terms of performance (Hoffmann et al., 2022; Wei et al., 2022b; Chowdhery et al., 2022), scale comes at a large cost in terms of hardware requirements and inference speed. Download statistics on HuggingFace show the smallest Llama model (Touvron et al., 2023a) to be the most adopted by the community, demonstrating the interest in small but capable models. LLM scaling laws demonstrate the diminishing returns of training a model of a given size past a certain amount of tokens. By continuing pre-training way past the compute-optimal threshold, performance has been shown not to saturate, enabling the training of “inference-optimal” models of high interest to the industrial and research communities (Sardana and Frankle, 2023). It is still not fully understood how model performance continues to improve at these later stages of training.
@@ 1.1 Contributions
Contribution 3: FrenchBench: A novel LLM benchmark for the French Language. To evaluate models in French, we introduce a benchmark encompassing various tasks to assess factual knowledge, generative capabilities, language understanding, etc. This benchmark is constructed both from openly available datasets, as well as newly released manually annotated data. We evaluate and report results for our models as well as other models with French-speaking capabilities. [^5] Code for evaluation is currently available at https://github.com/EleutherAI/lm-evaluation-harness [^6] Another complementary initiative has been led for French model evaluation and released concurrently in Bawden et al. (2024)
Contribution 4: Releasing high-performing, inference-optimal models for the industrial community, together with a wide range of resources for the research community. The models we train are all released under open licenses. Our largest model is trained on a 3000:1 token to parameter ratio (150 times longer than a Chinchilla Optimal 1.3B model) leading to very strong performance in its size category. We show that model performance on downstream tasks continues to dramatically improve with lengthened pre-training runs, although model perplexity does not significantly improve. We release all checkpoints for all model sizes, as well as the exact training data seen at every step for research purposes. [^7] Training data indexing order will be released in a second stage. These models are extremely efficient to run, leading to low-latency, energy-efficient on-edge inference, even on low-resource devices such as phones or personal computers. These releases [^8] Code for dataset collection and filtering is available at https://github.com/ManuelFay/llm-data-hub. Code for model training is hosted at https://github.com/CoderPat/croissant-llm-training. Datasets and model checkpoints are available at https://huggingface.co/CroissantLLM. are motivated by a commitment to transparency to allow research and reassuring users for industrial deployment: our models comply with 81 % of criteria listed by the Foundation Model Transparency Index (Bommasani et al., 2023) (see Section 6).
@@ 2.1 French Data
We ensure data is of good quality and correctly tagged in French through sampled manual inspection, and confirm French-speaking countries are well represented within the dataset. Notably, we include several news sources scrapped from Belgium, Switzerland, Canada, and Lebanon, as well as multiple African countries (Senegal, Morocco, Algeria, Cameroon, etc.)
@@
Encyclopedia Data. To introduce high-quality factual data to the training corpus, we integrate the French Wikipedia split from November 2023.212121https://huggingface.co/datasets/wikimedia/wikipedia with a CC-By-SA 3.0 license This corresponds to the latest knowledge cutoff in the training data. In total, more than 2.5 million articles are used, spanning more than 2 billion tokens.
Industrial Data. We scrap high-quality and publicly available data from industrial PDFs via a manually crafted list of websites, from large French and French Canadian (Quebec) companies to government agencies. This business-focused data boosts performance on a series of downstream applications related to industrial NLP. We collect over 36000 PDF multi-page documents and filter them through carefully crafted heuristics, followed by aggressive perplexity filtering. [^22] Data is public internet data that we release under MIT license with an opt-out scheme https://huggingface.co/datasets/manu/illuin_layout_dataset_text_only In total, we obtain over 290000 documents and 191 million tokens.
@@ 2.2 English Data
Internet Data. Similarly to the French dataset, we rely on carefully filtered content from an assortment of internet sources, including miscellaneous web pages and blogs. The filtering process includes heuristics and perplexity filtering, as well as large-scale deduplication (Soboleva et al., 2023). The SlimPajama corpus includes internet data from the CommonCrawl [^23] Common Crawl license https://commoncrawl.org/terms-of-use and C4242424ODC-BY license web scraps, as well as data sourced from Github textual content [^25] Only content under MIT, BSD, and Apache licenses are kept and the StackExchange forums. [^26] CC-By-SA 4.0 license https://archive.org/details/stackexchange
Miscellaneous. Other non-internet-based data sources are included in the SlimPajama dataset, such as scientific articles from Arxiv [^27] https://arxiv.org/, with author opt-out options and English documents from Wikipedia. [^28] CC-By-SA 3.0 license The SlimPajama dataset is also comprised of the “Books” subcorpus, obtained by downloading all book documents from Bibliotik. [^29] https://huggingface.co/datasets/the_pile_books3 Some of the documents within this last corpora have been flagged by their owner as proprietary data. We filter out all documents from this subcorpus, and replace them with data from the open-source Project Gutenberg (Hart, 1971) English books under public domains.303030 From https://huggingface.co/datasets/pg19 with an Apache 2.0 license
@@ 3 Training
Our main goal was to train a highly performing, yet resource-friendly bilingual model while optimizing performances across both languages. To focus on the specific challenges of the bilingual paradigm, we rely on previous work to motivate many of our design and hyperparameter choices (Touvron et al., 2023b).
@@ 3.2 Tokenizer
Most LLM tokenizers are fitted on English-centric corpora with an information-theoretic optimization objective, for example, Byte-Pair encoding (Sennrich et al., 2016) or Unigram (Kudo, 2018), leading to good fertility values (low token per word ratio) on English text, but high fertility in other languages. These phenomena make processing in other languages slower and more costly (Rust et al., 2021). Furthermore, subword splits in non-English languages mechanically carry less semantical meaning, potentially being a factor in the degraded performance of models on non-English languages (Rust et al., 2021).
@@ 3.3 Selecting an optimal language ratio
To offset this computational cost, we leverage recent findings on scaling laws (Kaplan et al., 2020a) that show that we can predict the performance of our model by training smaller models on the same dataset. In particular, Fernandes et al. (2023) found that, for multilingual models, by training smaller models with varying weights for each language in the data mix, one can fit a multilingual, joint scaling law that can predict the language performance trade-off of larger models, even for novel language weightings not encountered during the fitting of the scaling law.
@@
Figure 3 shows the performance predicted by jointly-fitted scaling laws as we scale the model and vary the language weightings on the Wiki data validation split. First, we see that the fitted scaling law is able to predict the performance of the larger model almost perfectly. Secondly, changing the weight of each language in training has a non-symmetrical impact on language performance: by increasing the (relative) weight of French from 50 % to 75 %, we get a marginal performance increase in French, while performance in English drops significantly. This fact is made clear by plotting the effective capacity ratio [^46] The number of parameters that it would take a monolingual model to match the performance of the multilingual, divided by the size of the multilingual model. See Fernandes et al. (2023) for more details of each language as we change the language weight (Figure 4): the “gains” in parameters from increasing weight of French data are minimal past the 50 % mark.
@@ 3.5 Training framework
We train our models on a modified version of Megatron-Deepspeed, [^47] https://github.com/deep-spin/Megatron-DeepSpeed a training framework built on top of PyTorch. Training is done on a dedicated Nvidia A100 SXM4 80 Gb partition of the Jean Zay supercomputer owned by the French government, on 30 octo-GPU nodes. Compute is obtained through a Jean Zay grant (2023-AD011014668R1). We rely on the HuggingFace Transformers and Datasets library for model and data manipulation.
To maximize efficiency, we set the micro-batch size per device to 8 sequences of length 2048, and use 4 gradient accumulation steps, resulting in a total batch size of $8\times 4\times 30\times 8=7680$ samples, or $7680*2048=15,728,640$ tokens. We achieve a mean efficiency of around 120 TFLOP484848160 TFLOP per second for our scaling law experiments with one GPU node only per second with activation checkpointing, leading to a total compute estimate of $224.30e$ 22 FLOPS. Standard Cross-Entropy losses are used on a Causal Language Modeling objective.
@@ 3.6 Training losses
Training lasts 17 days for a total of 99648 GPU hours, and we chose not to manually intervene, letting the model recover on its own after occasional loss spikes. We train with a max learning rate of $43e-$ 4, 1000 warmup steps, and a cosine learning rate with a minimum value of $51e-$ 5. Curves suggest the model still has not reached a performance plateau after 3T tokens (Figure 5). Checkpoints are stored every 5k steps and released with the rest of the project artifacts.
Refer to captionRefer to caption
@@ 3.7 Environmental impact
The model was exclusively trained on Jean Zay, a French government-owned supercomputer, operating on low-carbon nuclear electricity. Between experimental runs, scaling laws, and the final training, 123k A100 hours were used. The Thermal Design Power of the NVIDIA A100 SXM4 80Gb used is 400W corresponding to a total power consumption of 49.2 MWH and considering a grid carbon intensity of 57 gCO2eq/kWh, we estimate a carbon footprint of 2.80 tons of CO2 (Luccioni et al., 2022) emitted during training. [^49] Factoring in an intentionally large data center Power Usage Effectiveness of 1.2 (Luccioni et al., 2022), we estimate an emission of 3.36 tons of CO2.
Interestingly, the model we trained is not “compute-optimal” according to Chinchilla laws (Hoffmann et al., 2022), meaning that less computing could have been used to train a larger model with the same performance. However, our model aims to be used for inference purposes at industrial scales. Our training paradigm is thus to absorb the downstream inference costs, by training a smaller model on a lot more tokens to obtain an inference-optimized model equivalent in performance to a bigger compute-optimal model (Sardana and Frankle, 2023). Each inference of the final model is thus vastly more energy-efficient than a Chinchilla optimal model of equivalent performance ($>$ 3B parameters), and can even run on CPU or mobile devices. Relying on estimates of Kaplan et al. (2020b), at inference, CroissantLLM represents roughly 2.6 GFLOPS per token.
@@ 4 Evaluation Benchmarks
We hope to extend base model evaluation past English benchmarking alone and assess model capabilities in French, aiming for broad coverage across orthogonal capabilities to observe the effect of truly bilingual pre-training. Our evaluation efforts are rooted in transparency, and all results reported in the main technical report are reproducible through code that is open-sourced and public data. [^50] Evaluation is ran through the LM Evaluation harness (Gao et al., 2021) at https://github.com/EleutherAI/lm-evaluation-harness. Datasets used are grouped in the following collection: https://huggingface.co/collections/manu/frenchbench-evaluation-datasets-65bbcbf3987c4be8a5e9a2ef
@@ 5 Benchmark results
For monolingual French models, we use GPT-fr (Simoulin and Crabbé, 2021), a 1B model trained on 16.3B tokens, as well as the PagnolXL(1.5B) model (Launay et al., 2021), both in their author submitted HuggingFace implementations.
@@ 5.2.1 Chat Model
French Trivia. One main question this work attempts to tackle is whether training on bilingual data goes beyond augmenting the language understanding and writing capabilities of a model in another language, but also equips the models with novel knowledge and different cultural biases. We evaluate French cultural knowledge on a Trivia task, consisting of questions about France-related topics, asked in English (Table 6). As heuristic-based metrics are insufficient to capture the diversity of possible answers (Faysse et al., 2023), we rely on GPT-4 to score the results. Similar results to the MT-Bench experiments can be noted: although CroissantLLM is the best performing, no huge margins exist with respects to the best English model, TinyLlama, finetuned on the same bilingual SFT set. A potential bias might be that questions originate from American websites and are asked in English, which might slightly favor TinyLlama. Both CroissantLLM and TinyLlama largely outperform Bloom pretrained on a smaller corpora but finetuned on the same data. This further underlines the importance of pretraining data quantity, and the large knowledge overlap between various sources and languages.
Models Accuracy
Bloom 1b7 Chat 24.47 %
TinyLlamaChat 44.47 %
CroissantLLMChat 47.11 %
Table 6: French Trivia performance, evaluated by the gold answer and GPT4. Questions are asked in English and accuracy is measured by GPT-4. To remove bias due to Supervised Finetuning, all base models are finetuned on the same dataset.

### v5

# 1Illuin Technology  2Unbabel  3Equall  4Diabolocom  5MICS, CentraleSupélec 6Instituto de Telecomunicações, Lisboa  7INESC-ID, Lisboa  8Instituto Superior Técnico, Universidade de Lisboa 9Sorbonne Université, CNRS, ISIR, Paris    10Imperial College London

@@ 1 Introduction
Large Language Models (LLM) [^1] See Appendix E for a definition. have taken over the Natural Language Processing (NLP) landscape in the past years. Although a few proprietary models are still considered to run ahead of the pack (OpenAI et al., 2023), open weights models such as Llama (Touvron et al., 2023a; b), Qwen (Bai et al., 2023a) or Mistral (Jiang et al., 2023; 2024) are rapidly bridging the performance gap. However, widespread industrial and research adoption of such technology remains challenging for several reasons, including the lack of transparency in the data collection and training processes, the scarcity of existing resources outside of English, and the large-scale and costly nature of existing high-performing models.
@@
Challenging to use at scale. Although benefits of scaling models to enormous sizes have been amply demonstrated in terms of performance (Hoffmann et al., 2022; Wei et al., 2022b; Chowdhery et al., 2022), scale comes at a large cost in terms of hardware requirements and inference speed. Download statistics on HuggingFace show the smallest Llama model (Touvron et al., 2023a) to be the most adopted by the community, demonstrating the interest in small but capable models. LLM scaling laws demonstrate the diminishing returns of training a model of a given size past a certain amount of tokens. By continuing pre-training way past the compute-optimal threshold, performance has been shown not to saturate, enabling the training of “inference-optimal” models of high interest to the industrial and research communities (Sardana & Frankle, 2023). It is still not fully understood how model performance continues to improve at these later stages of training.
@@ 1.1 Contributions
Contribution 3: FrenchBench: A novel LLM benchmark for the French Language. To evaluate models in French, we introduce a benchmark encompassing various tasks to assess factual knowledge, generative capabilities, language understanding, etc. This benchmark is constructed both from openly available datasets, as well as newly released manually annotated data. We evaluate and report results for our models as well as other models with French-speaking capabilities. [^6] Code for evaluation is available at https://github.com/EleutherAI/lm-evaluation-harness [^7] Another complementary initiative has been led for French model evaluation and released concurrently in Bawden et al. (2024)
Contribution 4: Releasing high-performing, inference-optimal models for the industrial community, together with a wide range of resources for the research community. The models we train are all released under open licenses. Our largest model is trained on a 2300:1 token to parameter ratio (115 times longer than a Chinchilla Optimal 1.3B model) leading to very strong performance in its size category. We show that model performance on downstream tasks continues to dramatically improve with lengthened pre-training runs, although model perplexity does not significantly improve. We release all checkpoints for all model sizes, as well as the exact training data seen at every step for research purposes. [^8] Training data indexing order will be released in a second stage. These models are extremely efficient to run, leading to low-latency, energy-efficient on-edge inference, even on low-resource devices such as phones or personal computers. These releases [^9] Code for dataset collection and filtering is available at https://github.com/ManuelFay/llm-data-hub. Code for model training is hosted at https://github.com/CoderPat/croissant-llm-training. Datasets and model checkpoints are available at https://huggingface.co/CroissantLLM. are motivated by a commitment to transparency to allow research and reassuring users for industrial deployment: our models comply with 81 % of criteria listed by the Foundation Model Transparency Index (Bommasani et al., 2023) (see Section 6).
The CroissantLLM initiative aims to adress the aforementionned limitations of current models through the release of an inference-optimized, small but capable model that performs well outside of English settings, and that is designed to be as open and transparent as possible. Beyond facilitating industrial LLM adoption and unlocking new generative model use cases, this project is also a proven platform for researching LLMs and kickstarting future pretraining efforts (Martins et al., 2024; Meeus et al., 2024a; Shilov et al., 2024; Meeus et al., 2024b).
@@ 2.1 French Data
We ensure data is of good quality and correctly tagged in French through sampled manual inspection and confirm French-speaking countries are well represented within the dataset. Notably, we include several news sources scrapped from Belgium, Switzerland, Canada, and Lebanon, as well as multiple African countries (Senegal, Morocco, Algeria, Cameroon, etc.)  [^15] Contrary to certain languages such as Arabic or Portuguese in which the language differs greatly depending on the country, the written French language in these countries are largely similar with only a minor share of regionally introduced syntax and expressions
@@
Encyclopedia Data. To introduce high-quality factual data to the training corpus, we integrate the French Wikipedia split from November 2023.232323https://huggingface.co/datasets/wikimedia/wikipedia with a CC-By-SA 3.0 license This corresponds to the latest knowledge cutoff in the training data. In total, more than 2.5 million articles are used, spanning more than 2 billion tokens.
Industrial Data. We scrap high-quality and publicly available data from industrial PDFs via a manually crafted list of websites, from large French and French Canadian (Quebec) companies to government agencies. This business-focused data boosts performance on a series of downstream applications related to industrial NLP. We collect over 36000 PDF multi-page documents and filter them through carefully crafted heuristics, followed by aggressive perplexity filtering. [^24] Data is public internet data that we release under MIT license with an opt-out scheme https://huggingface.co/datasets/manu/illuin_layout_dataset_text_only. In total, we obtain over 290000 documents and 191 million tokens.
@@ 2.2 English Data
Internet Data. Similarly to the French dataset, we rely on carefully filtered content from an assortment of internet sources, including miscellaneous web pages and blogs. The filtering process includes heuristics and perplexity filtering, as well as large-scale deduplication (Soboleva et al., 2023). The SlimPajama corpus includes internet data from the CommonCrawl [^25] Common Crawl license https://commoncrawl.org/terms-of-use and C4262626ODC-BY license web scraps, as well as data sourced from Github textual content [^27] Only content under MIT, BSD, and Apache licenses are kept and the StackExchange forums. [^28] CC-By-SA 4.0 license https://archive.org/details/stackexchange
Miscellaneous. Other non-internet-based data sources are included in the SlimPajama dataset, such as scientific articles from Arxiv [^29] https://arxiv.org/, with author opt-out options and English documents from Wikipedia. [^30] CC-By-SA 3.0 license The SlimPajama dataset is also comprised of the “Books” subcorpus, obtained by downloading all book documents from Bibliotik. [^31] https://huggingface.co/datasets/the_pile_books3 Some of the documents within this last corpora have been flagged by their owner as proprietary data. We filter out all documents from this subcorpus, and replace them with data from the open-source Project Gutenberg (Hart, 1971) English books under public domains.323232 From https://huggingface.co/datasets/pg19 with an Apache 2.0 license
@@ 3 Training
Our main goal was to train a high-performing, yet resource-friendly bilingual model while optimizing performances across both languages. To focus on the specific challenges of the bilingual paradigm, we rely on previous work to motivate many of our design and hyperparameter choices (Touvron et al., 2023b).
@@ 3.2 Tokenizer
Most LLM tokenizers are fitted on English-centric corpora with an information-theoretic optimization objective, for example, Byte-Pair encoding (Sennrich et al., 2016) or Unigram (Kudo, 2018), leading to good fertility [^46] See Appendix E for a definition. values (low token per word ratio) on English text, but high fertility in other languages. These phenomena make processing in other languages slower and more costly (Rust et al., 2021). Furthermore, subword splits in non-English languages mechanically carry less semantical meaning, potentially being a factor in the degraded performance of models on non-English languages (Rust et al., 2021).
@@ 3.3 Selecting an optimal language ratio
To offset this computational cost, we leverage recent findings on scaling laws (Kaplan et al., 2020b) that show that we can predict the performance of our model by training smaller models on the same dataset. In particular, Fernandes et al. (2023) found that, for multilingual models, by training smaller models with varying weights for each language in the data mix, one can fit a multilingual, joint scaling law that can predict the language performance trade-off of larger models, even for novel language weightings not encountered during the fitting of the scaling law.
@@
Figure 3 shows the performance predicted by jointly-fitted scaling laws as we scale the model and vary the language weightings on the Wiki data validation split. First, we see that the fitted scaling law is able to predict the performance of the larger model almost perfectly. Secondly, changing the weight of each language in training has a non-symmetrical impact on language performance: by increasing the (relative) weight of French from 50 % to 75 %, we get a marginal performance increase in French, while performance in English drops significantly. This fact is made clear by plotting the effective capacity ratio [^49] The relative monolingual model size required to match the performance of the multilingual model on a language. See E for more details of each language as we change the language weight (Figure 4): the “gains” in parameters from increasing weight of French data are minimal past the 50 % mark.
@@ 3.5 Training framework
We train our models on a modified version of Megatron-Deepspeed, [^50] https://github.com/deep-spin/Megatron-DeepSpeed a training framework built on top of PyTorch. Training is done on a dedicated Nvidia A100 SXM4 80 Gb supercomputer partition with 30 octo-GPU nodes. We rely on the HuggingFace Transformers and Datasets library for model and data manipulation.
To maximize efficiency, we set the micro-batch size per device to 8 sequences of length 2048, and use 4 gradient accumulation steps, resulting in a total batch size of $8\times 4\times 30\times 8=7680$ samples, or $7680*2048=15,728,640$ tokens. We achieve a mean efficiency of around 120 TFLOP515151160 TFLOP per second for our scaling law experiments with one GPU node only per second with activation checkpointing, leading to a total compute estimate of $224.30e$ 22 FLOPS. Standard Cross-Entropy losses are used on a Causal Language Modeling objective.
@@ 3.6 Training losses
Training lasts 17 days for a total of 99648 GPU hours, and we chose not to manually intervene, letting the model recover on its own after occasional loss spikes.  [^52] Previous literature rarely mentions handling of such issues. Some models have reloaded training by skipping batches (Zhang et al., 2022; Anil et al., 2023), others have seemed to ignore it if loss recovered (Touvron et al., 2023a). In our case, we used gradient clipping, weight decay and tuned optimizer values (Zhang et al., 2022; Touvron et al., 2023b), to minimize the occurrences of such loss spikes… During training, as spikes were few and far between, and loss seemed to recover quite rapidly every time and without visible consequences, we opted to let the model train with no intervention. We highlight that since our checkpoints and dataloaders are openly available, this enables future experimentations around these loss spikes. We train with a max learning rate of $43e-$ 4, 1000 warmup steps, and a cosine learning rate with a minimum value of $51e-$ 5. Curves suggest the model still has not reached a performance plateau after 3T tokens (Figure 5). Checkpoints are stored every 5k steps and released with the rest of the project artifacts.
@@ 3.7 Environmental impact
The model was exclusively trained on a supercomputer operating on low-carbon nuclear electricity. Between experimental runs, scaling laws, and the final training, 123k A100 hours were used. The Thermal Design Power of the NVIDIA A100 SXM4 80Gb used is 400W corresponding to a total power consumption of 49.2 MWH and considering a grid carbon intensity of 57 gCO2eq/kWh, we estimate a carbon footprint of 2.80 tons of CO2 (Luccioni et al., 2022) emitted during training. [^53] Factoring in an intentionally large data center Power Usage Effectiveness of 1.2 (Luccioni et al., 2022), we estimate an emission of 3.36 tons of CO2.
Interestingly, the model we trained is not “compute-optimal” according to Chinchilla laws (Hoffmann et al., 2022), meaning that less computing could have been used to train a larger model with the same performance. However, our model aims to be used for inference purposes at industrial scales. Our training paradigm is thus to absorb the downstream inference costs, by training a smaller model on a lot more tokens to obtain an inference-optimized model equivalent in performance to a bigger compute-optimal model (Sardana & Frankle, 2023). Each inference of the final model is thus vastly more energy-efficient than a Chinchilla optimal model of equivalent performance ($>$ 3B parameters), and can even run on CPU or mobile devices. Relying on estimates of Kaplan et al. (2020a), at inference, CroissantLLM represents roughly 2.6 GFLOPS per token.
@@ 4 Evaluation Benchmarks
We hope to extend base model evaluation past English benchmarking alone and assess model capabilities in French, aiming for broad coverage across orthogonal capabilities to observe the effect of truly bilingual pre-training. Our evaluation efforts are rooted in transparency, and all results reported in the main technical report are reproducible through code that is open-sourced and public data. [^54] All benchmarks are made available and links will be provided in the non-anonymous version of this work.
@@ 5 Benchmark results
For monolingual French models, we use GPT-fr (Simoulin & Crabbé, 2021), a 1B model trained on 16.3B tokens, as well as the PagnolXL(1.5B) model (Launay et al., 2021), both in their author submitted HuggingFace implementations.
@@ 5.1 Base model
French Trivia. One main question this work attempts to tackle is whether training on bilingual data goes beyond augmenting the language understanding and writing capabilities of a model in another language, but also equips the models with novel knowledge and different cultural biases. We evaluate French cultural knowledge on a Trivia task, consisting of questions about France-related topics, asked in English (Table 5), and score results obtained in 5-shot settings with ROUGE-1.606060As heuristic-based metrics are often insufficient to capture the diversity of possible answers (Faysse et al., 2023), we also score predictions using GPT4 as a judge, and confirm the ROUGE1 metric is well suited for this task given the closed and short nature of the answers. CroissantLLM is the best performing model evaluated under the 7B size, outperforming English-centric models by significant margins. This knowledge gap showcases the effect of the pretraining data mix in specific knowledge acquisition, underlining the interest of integrating vast amounts of varied and multilingual data when training models aiming for broad knowledge coverage.